# PrismAudio: Decomposed Chain-of-Thoughts and Multi-dimensional Rewards for Video-to-Audio Generation

### v3

@@ Appendix A Related Work on Video-to-Audio Generation
V2A generation has recently gained significant attention with advances in multimodal AI systems (simeoni2025dinov3; brooks2024video; team2024gemini). Current V2A methods predominantly employ latent diffusion models (xing2024seeing; liu2025omniaudio; wang2024frieren), while some explore autoregressive token-based approaches. Diff-Foley (luo2023diff) and VTA-LDM (xu2024video) represent standard latent diffusion models conditioned on video features, while FoleyGen (mei2024foleygen) and V-AURA (viertola2025temporally) frame conditional audio generation as next-token prediction using visual features. Early methods focus on improving semantic consistency through better video representations. Works like CAVP (luo2023diff) and CLIP4CLIP (luo2022clip4clip) employ contrastive learning for video encoding, while adapter-based approaches like FoleyCrafter (zhang2024foleycrafter) build upon pre-trained text-to-audio models (liu2023audioldm; liu2024flashaudio; liu2024audiolcm) for enhanced controllability. Recent advances introduce explicit multimodal conditioning to address semantic limitations. MovieGen Audio (polyak2024movie) conditions on both text and video to generate video-aligned audio, achieving substantial progress and inspiring subsequent video-text-audio generation works (cheng2024taming; chen2025video; shan2025hunyuanvideo; tian2025audiox). Most recently, ThinkSound (liu2025thinksound) innovatively introduces CoT reasoning via MLLMs, replacing simple text prompts with structured reasoning that significantly improves interpretability and narrative coherence. However, existing V2A methods suffer from objective entanglement—optimizing competing perceptual goals through a single reconstruction loss—and lack human preference alignment beyond textual matching. In contrast, we propose PrismAudio, the first reinforcement learning framework for V2A generation with specialized multi-dimensional CoT-reward correspondence to address these fundamental limitations.
@@ B.1 From Deterministic ODEs to Stochastic SDEs
Generative modeling with flow matching (lipman2022flow) learns a velocity field $v_{\theta}(\mathbf{x}_{t},t,c)$ that transports a simple prior distribution (e.g., Gaussian noise) to a complex data distribution. The generation process is typically described by a deterministic probability flow Ordinary Differential Equation (ODE):
@@
Based on the principles of score-based generative modeling (song2020score), any such ODE has an equivalent Stochastic Differential Equation (SDE) that shares the same marginal probability distributions $p(\mathbf{x}_{t})$ at every time $t$. For a rectified flow backbone, the velocity field $v_{\theta}$ is an approximation of the drift term. We can construct the corresponding SDE by re-deriving the drift and adding a diffusion term. Specifically, the full SDE can be written as:
@@ B.3 Derivation of the Per-Step Policy and Ratio
which is the closed-form computation of the GRPO policy ratio (Eq. 4) used in our final objective function (Eq. 7). Following the practice of  liu2025flow, we use KL regularization to mitigate reward hacking. The validation of KL regularization is presented in Appendix E.5.
@@ C.1 Benchmark Construction
To create AudioCanvas, we target sound effects and music as primary audio categories, drawing inspiration from the AudioSet ontology (gemmeke2017audio). We first refine AudioSet by filtering out categories related to human speech and singing, as well as rare classes with insufficient data. This results in a target of 300 distinct categories relevant to V2A generation.
Our construction process involves two main stages. In the first stage, we analyze existing benchmarks like Kling-Audio-Eval (wang2025kling) and find that it covers only a fraction of our target classes ( 100 out of 300) and often features overly simple scenarios. To address this limitation, we develop a rigorous filtering protocol. We start with a large pool of candidate videos and automatically filter out samples where existing V2A models already achieve near-perfect scores (low FD and KL scores, see Section 4 for evaluation metrics), as these scenarios pose no new challenge. The remaining samples are then manually screened by professional audio experts to exclude videos with minimal diversity, repetitive sounds, or simple visual contexts. This multi-step process ensures that only high-quality, challenging samples are retained.
@@ C.2 Quality Assessment and Benchmark Comparison
Clotho (drossos2020clotho) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
AudioCaps (kim2019audiocaps) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
Epic Sounds (huh2025epic) Action2Sound 10K 44 Video+Audio High (Egocentric) Class Label Unspecified
AudioSet (gemmeke2017audio) Classification 18K 527 Video+Audio Low (Automatic) Class Label Unspecified
VGGSound (chen2020vggsound) Classification & V2A 15K 309 Video+Audio Low (Known Issues) Class Label Unspecified
Kling-Audio-Eval (wang2025kling) V2A 21K 100 Video+Audio Moderate (Curated) Simple Caption Primarily Single-event
@@ Appendix D Implementation Details
For VAE training, we fine-tune our variational autoencoder on stereo audio data at 44.1kHz sample rate using the foundation provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools, employing mixed precision training with a batch size of 144 distributed across 24 A800 GPUs for 500,000 training steps. For the audio foundation model, we integrate VideoPrism-Large  [^4] https://huggingface.co/google/videoprism-lvt-large-f8r288 as the video encoder and T5-Gemma Large  [^5] https://huggingface.co/google/t5gemma-l-l-ul2-it as the text encoder. The pre-training phase of the audio foundation model uses WavCaps (Mei_2024), AudioCaps (kim2019audiocaps), and VGGSound (chen2020vggsound) datasets. We utilize exponential moving average (EMA) and automatic mixed precision (AMP) for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt classifier-free guidance (CFG) (ho2022classifier), dropout of 0.1 for each modality with a learning rate of 1e-4. For Chain-of-Thought fine-tuning, we continue training the pre-trained model on our curated multi-dimensional CoT dataset, which is annotated using the VGGSound dataset by the fine-tuned VideoLLaMA2, using the same configuration of hyperparameters. For the reinforcement learning post-training using the VGGSound dataset, we fine-tune the audio foundation model with a learning rate of 1e-5. The Fast-GRPO hyperparameters are configured as follows: KL ratio of 0.04, noise level of 0.7, group size of 16, SDE steps of 2, and sampling steps of 24.
@@ E.2 Multi-dimensional CoT and RL Analysis on VGGSound
Multi-dimensional vs. Single-dimensional Rewards. To supplement our main analysis on AudioCanvas, we conduct the same reward ablation study on the in-domain VGGSound test set. The results, presented in Table 9, consistently validate our core findings: (1) Single-dimensional optimization causes severe objective entanglement. Echoing the findings on AudioCanvas, optimizing for a single goal leads to catastrophic imbalances. For instance, Semantic Only optimization boosts the CLAP score (0.51) but severely degrades temporal synchrony (DeSync: 0.66). The most dramatic example is the Aesthetic Only model, whose high PQ score (6.98) comes at the cost of a nearly quadrupled distribution error (FD: 1.10 $\rightarrow$ 4.27), confirming that this approach produces pleasing but content-detached audio. (2) Our multi-dimensional reward successfully balances competing objectives. In stark contrast, our full Multi-dimensional approach once again demonstrates its ability to navigate and balance these inherent objective tensions. It is the only method that achieves holistic improvement, simultaneously enhancing semantics (CLAP: 0.44 $\rightarrow$ 0.47), temporal synchrony (DeSync: 0.48 $\rightarrow$ 0.41), and aesthetic quality (PQ: 6.24 $\rightarrow$ 6.38) over the baseline. This confirms that the necessity for holistic optimization is a fundamental principle, holding true across different data distributions.
@@ E.5 Analysis of Reward Hacking Mitigation
A common challenge in reinforcement learning is “reward hacking," where a model exploits a reward proxy without achieving the intended goal. To mitigate this, we adopt the practice of Flow-GRPO liu2025flow by incorporating a KL penalty with a weight of 0.04 into our objective function. This regularizes the policy update, preventing it from deviating too drastically from the stable pre-trained model, thus discouraging "exploitative" shortcuts to high rewards.
@@ E.6 Empirical Validation of ODE-SDE Distribution Equivalence
Figure 5: Empirical validation of ODE–SDE distribution equivalence during training.
@@ E.7 Ablation Study on Aesthetic and Spatial Rewards
To rigorously validate the necessity of our full four-dimensional framework, we conduct comprehensive ablation studies on the VGGSound test set, evaluating three reduced configurations: (1) removing Aesthetic reward (Sem+Temp+Spatial), (2) removing Spatial reward (Sem+Temp+Aesthetic), and (3) removing both (Sem+Temp only). The results are presented in Table 13.
Quantitative Analysis:
(1) Sem+Temp only configuration: While achieving improvements in semantic (CLAP: 0.44 $\rightarrow$ 0.48) and temporal (DeSync: 0.48 $\rightarrow$ 0.41) metrics, this configuration shows no improvement in spatial accuracy (CRW: 8.29 $\rightarrow$ 8.38, still far from optimal) and only marginal aesthetic gains (PQ: 6.24 $\rightarrow$ 6.30). More critically, the Frechet Distance increases (1.10 $\rightarrow$ 1.22), indicating lower overall audio quality. This demonstrates that aesthetic and spatial qualities are not "free" benefits from semantic/temporal optimization.
(2) Sem+Temp+Spatial configuration: Adding spatial reward improves spatial metrics (CRW: 8.38 $\rightarrow$ 7.53, GCC: 4.11 $\rightarrow$ 3.75) and distribution quality (FD: 1.22 $\rightarrow$ 1.03), but fails to enhance aesthetics (PQ: 6.30 $\rightarrow$ 6.19, actually degrades). This confirms that aesthetic quality requires explicit optimization.
(3) Sem+Temp+Aesthetic configuration: Adding aesthetic reward improves aesthetic metrics (PQ: 6.30 $\rightarrow$ 6.44, PC: 3.25 $\rightarrow$ 3.11, CE: 3.98 $\rightarrow$ 4.17, CU: 5.61 $\rightarrow$ 5.76), but degrades distribution quality (FD: 1.22 $\rightarrow$ 1.52). This confirms that spatial accuracy requires explicit optimization.
(4) Full four-dimensional approach: This is the only configuration achieving holistic improvement across all axes simultaneously: semantic (CLAP: 0.44 $\rightarrow$ 0.47), temporal (DeSync: 0.48 $\rightarrow$ 0.41), aesthetic (PQ: 6.24 $\rightarrow$ 6.38, CE: 3.94 $\rightarrow$ 4.29), spatial (CRW: 8.29 $\rightarrow$ 7.72), and distribution quality (FD: 1.10 $\rightarrow$ 1.08, KL: 1.24 $\rightarrow$ 1.23). This confirms that a multi-dimensional reward system is essential to disentangle competing objectives and prevent "reward hacking."

### v4

@@ 5 Conclusion
Reproducibility Statement
To ensure reproducibility, the code, the AudioCanvas benchmark, and all model weights will be made publicly available upon publication. Core implementation details are provided in Appendix D. The released package will include:
  • Complete training scripts and configuration files required to reproduce the main results.
  • The training dataset generated by Gemini 2.5 Pro for the VideoLLaMA2 fine-tuning stage.
  • Detailed documentation covering the experimental setup and hyperparameter settings.
Acknowledgement
The research was supported in part by Early Career Scheme (ECS-HKUST22201322), Theme-based Research Scheme under Grant T45-205/21-N from Hong Kong RGC, and Generative AI Research and Development Centre from InnoHK.
@@ Appendix A Related Work on Video-to-Audio Generation
V2A generation has recently gained significant attention with advances in multimodal AI systems (Siméoni et al., 2025; Brooks et al., 2024; Team et al., 2024a). Current V2A methods predominantly employ latent diffusion models (Xing et al., 2024; Liu et al., 2025a; Wang et al., 2024b), while some explore autoregressive token-based approaches. Diff-Foley (Luo et al., 2023) and VTA-LDM (Xu et al., 2024) represent standard latent diffusion models conditioned on video features, while FoleyGen (Mei et al., 2024b) and V-AURA (Viertola et al., 2025) frame conditional audio generation as next-token prediction using visual features. Early methods focus on improving semantic consistency through better video representations. Works like CAVP (Luo et al., 2023) and CLIP4CLIP (Luo et al., 2022) employ contrastive learning for video encoding, while adapter-based approaches like FoleyCrafter (Zhang et al., 2024) build upon pre-trained text-to-audio models (Liu et al., 2023a; 2024b; 2024a) for enhanced controllability. Recent advances introduce explicit multimodal conditioning to address semantic limitations. MovieGen Audio (Polyak et al., 2024) conditions on both text and video to generate video-aligned audio, achieving substantial progress and inspiring subsequent video-text-audio generation works (Cheng et al., 2024a; Chen et al., 2025; Shan et al., 2025; Tian et al., 2025). Most recently, ThinkSound (Liu et al., 2025b) innovatively introduces CoT reasoning via MLLMs, replacing simple text prompts with structured reasoning that significantly improves interpretability and narrative coherence. However, existing V2A methods suffer from objective entanglement—optimizing competing perceptual goals through a single reconstruction loss—and lack human preference alignment beyond textual matching. In contrast, we propose PrismAudio, the first reinforcement learning framework for V2A generation with specialized multi-dimensional CoT-reward correspondence to address these fundamental limitations.
@@ B.1 From Deterministic ODEs to Stochastic SDEs
Generative modeling with flow matching (Lipman et al., 2022) learns a velocity field $v_{\theta}(\mathbf{x}_{t},t,c)$ that transports a simple prior distribution (e.g., Gaussian noise) to a complex data distribution. The generation process is typically described by a deterministic probability flow Ordinary Differential Equation (ODE):
@@
Based on the principles of score-based generative modeling (Song et al., 2020), any such ODE has an equivalent Stochastic Differential Equation (SDE) that shares the same marginal probability distributions $p(\mathbf{x}_{t})$ at every time $t$. For a rectified flow backbone, the velocity field $v_{\theta}$ is an approximation of the drift term. We can construct the corresponding SDE by re-deriving the drift and adding a diffusion term. Specifically, the full SDE can be written as:
@@ B.3 Derivation of the Per-Step Policy and Ratio
which is the closed-form computation of the GRPO policy ratio (Eq. 4) used in our final objective function (Eq. 7). Following the practice of  Liu et al. (2025c), we use KL regularization to mitigate reward hacking. The validation of KL regularization is presented in Appendix E.5.
@@ C.1 Benchmark Construction
To create AudioCanvas, we target sound effects and music as primary audio categories, drawing inspiration from the AudioSet ontology (Gemmeke et al., 2017). We first refine AudioSet by filtering out categories related to human speech and singing, as well as rare classes with insufficient data. This results in a target of 300 distinct categories relevant to V2A generation.
Our construction process involves two main stages. In the first stage, we analyze existing benchmarks like Kling-Audio-Eval (Wang et al., 2025) and find that it covers only a fraction of our target classes ( 100 out of 300) and often features overly simple scenarios. To address this limitation, we develop a rigorous filtering protocol. We start with a large pool of candidate videos and automatically filter out samples where existing V2A models already achieve near-perfect scores (low FD and KL scores, see Section 4 for evaluation metrics), as these scenarios pose no new challenge. The remaining samples are then manually screened by professional audio experts to exclude videos with minimal diversity, repetitive sounds, or simple visual contexts. This multi-step process ensures that only high-quality, challenging samples are retained.
@@ C.2 Quality Assessment and Benchmark Comparison
Clotho (Drossos et al., 2020) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
AudioCaps (Kim et al., 2019) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
Epic Sounds (Huh et al., 2025) Action2Sound 10K 44 Video+Audio High (Egocentric) Class Label Unspecified
AudioSet (Gemmeke et al., 2017) Classification 18K 527 Video+Audio Low (Automatic) Class Label Unspecified
VGGSound (Chen et al., 2020) Classification & V2A 15K 309 Video+Audio Low (Known Issues) Class Label Unspecified
Kling-Audio-Eval (Wang et al., 2025) V2A 21K 100 Video+Audio Moderate (Curated) Simple Caption Primarily Single-event
@@ Appendix D Implementation Details
For VAE training, we fine-tune our variational autoencoder on stereo audio data at 44.1kHz sample rate using the foundation provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools, employing mixed precision training with a batch size of 144 distributed across 24 A800 GPUs for 500,000 training steps. For the audio foundation model, we integrate VideoPrism-Large  [^4] https://huggingface.co/google/videoprism-lvt-large-f8r288 as the video encoder and T5-Gemma Large  [^5] https://huggingface.co/google/t5gemma-l-l-ul2-it as the text encoder. The pre-training phase of the audio foundation model uses WavCaps (Mei et al., 2024a), AudioCaps (Kim et al., 2019), and VGGSound (Chen et al., 2020) datasets. We utilize exponential moving average (EMA) and automatic mixed precision (AMP) for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt classifier-free guidance (CFG) (Ho and Salimans, 2022), dropout of 0.1 for each modality with a learning rate of 1e-4. For Chain-of-Thought fine-tuning, we continue training the pre-trained model on our curated multi-dimensional CoT dataset, which is annotated using the VGGSound dataset by the fine-tuned VideoLLaMA2, using the same configuration of hyperparameters. For the reinforcement learning post-training using the VGGSound dataset, we fine-tune the audio foundation model with a learning rate of 1e-5. The Fast-GRPO hyperparameters are configured as follows: KL ratio of 0.04, noise level of 0.7, group size of 16, SDE steps of 2, and sampling steps of 24.
D.1 GPU Resource Requirements
Inference Requirements.
The proposed PrismAudio requires modest GPU resources for inference. When running on NVIDIA A800 GPUs with batchsize=1 (single sample generation), the inference process consumes approximately 5,618 MiB of VRAM.
Training Requirements.
The training pipeline consists of several stages:
  1. 1.
    VAE Fine-tuning (Optional): Requires 24 GPUs (NVIDIA 80GB A800) for approximately 5 days. This is the most computationally intensive stage. However, VAE fine-tuning is optional and can be skipped with a performance trade-off (as discussed in our response to Reviewer BAuq). We will provide pre-trained VAE checkpoints that can be used directly without fine-tuning.
  2. 2.
    Main Model Training (Flow Matching): Requires 16 GPUs (NVIDIA 80GB A800) for approximately 3 days. This stage trains the audio foundation model using flow matching on the pre-training datasets (WavCaps, AudioCaps, VGGSound).
  3. 3.
    Fast-GRPO Post Training: Requires 8 GPUs (NVIDIA 80GB A800) for approximately 5 days. The windowed SDE sampling significantly reduces computational cost compared to full SDE-GRPO (approximately 8 days with the same GPU resources), achieving a 1.6× speedup.
  4. 4.
    VideoLLaMA2 Fine-tuning: Requires 8 GPUs (NVIDIA 80GB A800) for approximately 2 days. This stage fine-tunes VideoLLaMA2 to generate multi-dimensional CoT descriptions from video inputs.
Total Training Cost.
The complete training pipeline requires approximately 16-24 GPUs over 2-3 weeks, depending on the specific configuration. We acknowledge that the training phase requires substantial computational resources, but several important points need to be noted:
  • The VAE fine-tuning stage is optional and can be skipped with a performance trade-off.
  • Fast-GRPO training is significantly more efficient than full SDE-GRPO, reducing training time by approximately 1.6×.
  • We will release all pre-trained checkpoints so researchers can directly use the model for inference without retraining.
D.2 CoT Generation Prompts
Prompt for Gemini 2.5 Pro (AudioCanvas and VideoLLaMA2 Training Data).
We use the following prompt to instruct Gemini 2.5 Pro to generate structured CoT descriptions covering four dimensions: semantic, temporal, aesthetic, and spatial. This prompt is used both for constructing the AudioCanvas benchmark and for generating training data for VideoLLaMA2 fine-tuning:
You are an expert in video-to-audio generation. Given a video with audio, analyze the audio content that should be generated and provide a comprehensive Chain-of-Thought description covering four dimensions: Semantic Dimension: Identify all audio events, objects, and actions visible in the video. Describe what sounds should be generated, including their characteristics (e.g., type, intensity, material properties). Temporal Dimension: Determine the sequential ordering and timing of audio events. Describe when each sound should occur relative to visual cues and other audio events, including onset, duration, and temporal relationships. Aesthetic Dimension: Assess the audio quality aspects that should be achieved. Describe the naturalness, fidelity, richness, and perceptual quality of the sounds, considering the context and environment depicted in the video. Spatial Dimension: Analyze the spatial positioning of sound sources. Describe the directional placement, left-right channel distribution, distance, and movement patterns of sounds relative to the visual content. Provide your analysis in a structured format that clearly separates these four dimensions.
Prompt for Text LLM (Multi-dimensional CoT Transformation).
To transform the CoT captions generated by Gemini 2.5 Pro into our desired multi-dimensional decoupled CoT input format, we use a text LLM with the following prompt:
Transform the following Chain-of-Thought description into four separate, decoupled CoT modules. Extract and reorganize the content into: Semantic CoT: Extract only the semantic content (audio events, objects, actions, characteristics). Format as a focused reasoning text for semantic audio generation. Temporal CoT: Extract only the temporal content (sequencing, timing, onset, duration, temporal relationships). Format as a focused reasoning text for temporal synchronization. Aesthetic CoT: Extract only the aesthetic content (naturalness, fidelity, quality, perceptual aspects). Format as a focused reasoning text for aesthetic audio generation. Spatial CoT: Extract only the spatial content (directional placement, left-right distribution, distance, movement). Format as a focused reasoning text for spatial audio generation. Ensure each CoT module is self-contained and focused solely on its respective dimension, without cross-dimensional references.
D.3 VideoLLaMA2 Fine-tuning Details
We use the official VideoLLaMA2 repository’s fine-tuning code with DeepSpeed ZeRO-3. We initialize from the pre-trained VideoLLaMA2-AV (7B) model and employ AdamW optimizer with learning rate $2\times 10^{-5}$ and weight decay $0.0$, using cosine annealing with warmup (warmup ratio $0.03$). The training uses a batch size of 4 per GPU with global batch size 128 (via gradient accumulation, which is automatically calculated based on world size and number of GPUs), and runs for 10 epochs with standard next-token prediction loss. Frozen components: Video encoder, audio encoder, and audio projector are kept frozen. Trainable components: Only the video projector and language model (LLM) are updated during fine-tuning. By completely freezing the audio components, we force the model to learn better visual representations to compensate for the absence of audio information in silent videos. This design helps bridge the gap between training (where we use sounding videos to generate CoTs) and inference (where we use silent videos), ensuring that the model learns to generate high-quality CoTs from visual information alone.
@@ E.2 Multi-dimensional CoT and RL Analysis on VGGSound
Multi-dimensional vs. Single-dimensional Rewards. To supplement our main analysis on AudioCanvas, we conduct the same reward ablation study on the in-domain VGGSound test set. The results, presented in Table 9, consistently validate our core findings: (1) Single-dimensional optimization causes severe objective entanglement. Echoing the findings on AudioCanvas, optimizing for a single goal leads to catastrophic imbalances. For instance, Semantic Only optimization boosts the CLAP score (0.51) but severely degrades temporal synchrony (DeSync: 0.66). The most dramatic example is the Aesthetic Only model, whose high PQ score (6.98) comes at the cost of a nearly quadrupled distribution error (FD: 1.14 $\rightarrow$ 4.27), confirming that this approach produces pleasing but content-detached audio. (2) Our multi-dimensional reward successfully balances competing objectives. In stark contrast, our full Multi-dimensional approach once again demonstrates its ability to navigate and balance these inherent objective tensions. It is the only method that achieves holistic improvement, simultaneously enhancing semantics (CLAP: 0.44 $\rightarrow$ 0.47), temporal synchrony (DeSync: 0.48 $\rightarrow$ 0.41), and aesthetic quality (PQ: 6.17 $\rightarrow$ 6.38) over the baseline. This confirms that the necessity for holistic optimization is a fundamental principle, holding true across different data distributions.
@@ E.5 Analysis of Reward Hacking Mitigation
A common challenge in reinforcement learning is “reward hacking,” where a model exploits a reward proxy without achieving the intended goal. To mitigate this, we adopt the practice of Flow-GRPO Liu et al. (2025c) by incorporating a KL penalty with a weight of 0.04 into our objective function. This regularizes the policy update, preventing it from deviating too drastically from the stable pre-trained model, thus discouraging ”exploitative” shortcuts to high rewards.
@@ E.6 Empirical Validation of ODE-SDE Distribution Equivalence
Figure 5: Empirical validation of ODE–SDE distribution equivalence during training. The figure tracks representative samples across training steps, where star markers represent ODE-generated results and circles of the same color represent SDE-generated results from the same input. The close alignment of ODE and SDE distributions throughout training demonstrates that mixed ODE–SDE sampling preserves the terminal distribution with high fidelity in practice, even under iterative parameter updates.
@@ E.7 Ablation Study on Aesthetic and Spatial Rewards
To rigorously validate the necessity of our full four-dimensional framework, we conduct comprehensive ablation studies on the VGGSound test set, evaluating three reduced configurations: (1) removing Aesthetic reward (Sem+Temp+Spatial), (2) removing Spatial reward (Sem+Temp+Aesthetic), and (3) removing both (Sem+Temp only). The results are presented in Table 13. Quantitative Analysis: (1) Sem+Temp only configuration: While achieving improvements in semantic (CLAP: 0.44 $\rightarrow$ 0.48) and temporal (DeSync: 0.48 $\rightarrow$ 0.41) metrics, this configuration shows no improvement in spatial accuracy (CRW: 8.29 $\rightarrow$ 8.38, still far from optimal) and only marginal aesthetic gains (PQ: 6.24 $\rightarrow$ 6.30). More critically, the Frechet Distance increases (1.10 $\rightarrow$ 1.22), indicating lower overall audio quality. This demonstrates that aesthetic and spatial qualities are not ”free” benefits from semantic/temporal optimization. (2) Sem+Temp+Spatial configuration: Adding spatial reward improves spatial metrics (CRW: 8.38 $\rightarrow$ 7.53, GCC: 4.11 $\rightarrow$ 3.75) and distribution quality (FD: 1.22 $\rightarrow$ 1.03), but fails to enhance aesthetics (PQ: 6.30 $\rightarrow$ 6.19, actually degrades). This confirms that aesthetic quality requires explicit optimization. (3) Sem+Temp+Aesthetic configuration: Adding aesthetic reward improves aesthetic metrics (PQ: 6.30 $\rightarrow$ 6.44, PC: 3.25 $\rightarrow$ 3.11, CE: 3.98 $\rightarrow$ 4.17, CU: 5.61 $\rightarrow$ 5.76), but degrades distribution quality (FD: 1.22 $\rightarrow$ 1.52). This confirms that spatial accuracy requires explicit optimization. (4) Full four-dimensional approach: This is the only configuration achieving holistic improvement across all axes simultaneously: semantic (CLAP: 0.44 $\rightarrow$ 0.47), temporal (DeSync: 0.48 $\rightarrow$ 0.41), aesthetic (PQ: 6.24 $\rightarrow$ 6.38, CE: 3.94 $\rightarrow$ 4.29), spatial (CRW: 8.29 $\rightarrow$ 7.72), and distribution quality (FD: 1.10 $\rightarrow$ 1.08, KL: 1.24 $\rightarrow$ 1.23). This confirms that a multi-dimensional reward system is essential to disentangle competing objectives and prevent ”reward hacking.”